Calculate the number of N atoms in 33.0 mg of NH4NO2.

 

1
Expert's answer
2022-03-13T15:18:39-0400

molar mass of NH4NO2 = 64.0 g/mol


moles of NH4NO2 = mass / molar mass = (92.6/1000) g / 64.0 g/mol = 0.00145 mol


1 mole of NH4NO2 contains two moles of N atoms,


thus, moles of N atoms = moles of NH4NO2 x 2 = 0.00145 x 2 = 0.00290 mol


1 mole of N atoms contains 6.02 x 10^23 N atoms,


number of N atoms = 0.00290 x 6.02 x 10^23 = 1.75 x 10^21